# Unmasking Social Bots: How Confident Are We?

### v1

@@ Abstract
Introduction
@@
Social media platforms have fundamentally transformed global communication, enabling the near-instant dissemination of information. Platforms like Facebook, YouTube, and Twitter/X, have over two billion monthly active users [1] and empower individuals to broadcast their thoughts easily. However, the popularity that social media enjoys has incentivized malicious actors such as tech-savvy individuals or governments [2], to deploy social bots to influence organic discourse and manipulate social media users for economic or political profit. Social bots [3, 4] — accounts controlled partly or fully by software — have been used to artificially boost the popularity of political candidates [5], spread conspiracy theories [6, 7] during health crises [8], and to manipulate the stock market [9, 10].
@@
A 2023 study that analyzed 1 million tweets during a Republican debate and a Donald Trump interview identified over 1,200 bot accounts spreading false narratives, highlighting the significant impact of bots during major events [11]. Anecdotal reports further underscore the prevalence of bot accounts, with some users suggesting that a substantial fraction of interactions on the platform are bot-generated. Despite enhanced verification processes, verified accounts continue to promote fraudulent schemes, illustrating the ongoing challenge in curbing bot activity.
@@
The issue of bot prevalence has also featured prominently in high-profile disputes, such as Elon Musk’s acquisition of Twitter. Musk’s legal team, using Botometer [12], an online tool for identifying spam and fake accounts, claimed that 33% of “visible accounts” were “false or spam”. However, Botometer’s creator, Kaicheng Yang, criticized this claim and their methodology, stating that the figure was misleading and disclosed that Musk’s team had not consulted him before using the tool [13]. Perhaps even more concerning is the adaptation of AI to bypass security systems deemed “prove you’re not a robot tests.” This concern has been echoed by Musk, Fig. 1 (left), as these tests are commonly used as initial filters for the prevention of inauthentic and/or malicious bot accounts. Moreover, accounts that successfully bypass such tests are increasingly becoming more “human-like,” as indicated by the dimensionality reduced representations of accounts in Fig. 1 (right). Musk has also implicated that the only reasonable method to solve the bot crisis is to force Twitter users into a paid subscription. Unfortunately, paid subscription has only worsened the bot-problem, since fake accounts can now pay to receive verification status (blue check mark), and thus appear authentic, as highlighted by the European Commission claim that Twitter is in violation the Digital Services Act [14]. Similarly, Elon Musk’s imposition of a prohibitively costly paywall to access Twitter’s research API has hampered the ability of researchers to study and mitigate bot activities.
@@
Figure 1: Impact of Bot Crisis and Uncertainty on Predictions (Our Work): As stated by Musk through X (left), the ability of bots to replicate human behavior and bypass security measures has increased dramatically with the advent of AI. Bot accounts are able to more efficiently mask themselves within the human population on social media platforms. This is shown through dimensionally reduced representations (right), in which we show three planes: (i) the true distributions, where we introduce an offset between human and bot points to ease visualization; (ii) the predictions of our network along with (iii) the associated uncertainty across the feature space. For (ii) and (iii) we use Gaussian process regression [15] for visualization purposes. Uncertainty is greater in regions where ambiguity is higher and the two classes overlap.
@@
Given the serious harm that social bots wielded by bad actors, pose to democracy [16, 17], public health [18, 19, 20], or the the economy [21], researchers have responded by developing a broad range of bot detection tools. Various Machine-Learning methods have been trained to detect social bots with a combination of features extracted from the social network structure, content/profile attributes, and temporal patterns [3]. Alternatively, all of these feature types are combined into a single model [22, 23, 24, 25, 26, 27]. Bot-detection algorithms often start by modeling the characteristics of accounts as the first step to distinguish bot from human-like behaviors [4]. Accounts may be represented using user profile information [28], content [29, 30, 31], actions [32], social network [33], and temporal signatures [32].
@@
Despite the progress made, bot detection remains a challenging, unsolved problem that is fraught with uncertainty due to the heterogeneity of bot behaviors, training data, and detection algorithms [34]. Consequently, it is not surprising for detection models to disagree on whether to label the same account as bot or human, since they are often trained on different datasets that are sensitive to different subsets of the signals of automation. In light of this, it is paramount that bot detection models provide uncertainty weights alongside the probability estimate that an account is bot-controlled, as this could determine if one should trust a prediction. However, existing bot detection algorithms focus exclusively on detection.
@@
Our work addresses both bot detection and the quantification of uncertainty at the account level, an important distinctive feature of this research. This dual focus is crucial as it allows us to leverage additional information related to the quantified uncertainty of each prediction, thereby enhancing decision-making and improving the reliability of bot classifications. Specifically, our approach facilitates targeted interventions for bots where predictions are made with high confidence and suggests a cautious approach, e.g., gathering more data for bots where predictions are uncertain. Also, our method is agnostic to bot-detection algorithm, which enables the inclusion of uncertainty measurement into existing bot detection systems. As discussed earlier, research at the nexus of uncertainty quantification and bot detection is both timely and novel. In subsequent sections, we present our methodology that adeptly separates aleatoric and epistemic uncertainties. Aleatoric uncertainty, arising from the inherent randomness in our bot datasets, is characterized using multiple features from the BLOC framework [35], or features from Botometer [12]. Conversely, epistemic uncertainty originates from the limitations of the predictive model used. Previous studies such as [36] have employed Deep Ensemble and Stochastic Weight Averaging (SWA) for uncertainty estimation, yet these methods have notable limitations [37]. SWA does not differentiate between the two types of uncertainties and requires supplementary techniques like dropout or bootstrapping for effective uncertainty quantification. Similarly, while Deep Ensembles capture epistemic uncertainty, they fail to distinctly separate it from aleatoric uncertainty.
@@
In contrast, our approach utilizes Bayesian methods, which provide a comprehensive and theoretically grounded framework for distinguishing and quantifying different types of uncertainties. We demonstrate that the computational demand of our Bayesian neural network is negligible given the complexity of the problem at hand, and agnostic to input, e.g. BLOC or Botometer features. This approach positions our methodology as exceptionally capable in the critical domain of bot detection, offering a robust mechanism for managing uncertainty quantification.
Results
In this section, we discuss the results of predicting the labels (bot or human) of accounts with our network, a Bayesian Neural Network (BNN) inspired by the Event-Level-Uncertainty Quantification (ELUQuant) [38] work originally developed for nuclear physics. This approach allows for the calculation of aleatoric and epistemic uncertainty in the predictions of whether Twitter/X accounts are bots or humans.
It is also important to compare the performance of the BNN with its deterministic counterpart, the Deep Neural Network (DNN). A BNN should perform at least as well as its deterministic counterpart. This has been demonstrated in the following way. We evaluate our model using standard methods, namely the Receiver Operating Characteristic (ROC) Curve and the associated Area Under the Curve (AUC). These metrics allow us to avoid making hard threshold cuts in the probability space and reflect a model’s performance across various thresholds. Thus, AUC is an ideal metric for such use cases. We then compare the results to the deterministic counterpart of our BNN, a DNN, along with the Random Forest (RF) from Nwala et al. [35]. We extract both BLOC and Botometer features for the same set of users. Figure 2 shows a comparison between three methods, with the AUC indicated in the legend. The left plot contains the ROC curves for the algorithms trained on BLOC features, where the error is calculated through bootstrapping over the posterior on the probability. This approach allows us to produce $5\sigma$ bands over both the True Positive Rate (TPR) and False Positive Rate (FPR). The right plot provides the same ROC Curves for algorithms trained on Botometer features. We note an AUC for the BNN of $0.966\pm 0.001$, which agrees with the deterministic DNN that achieves an AUC of $0.969$ on BLOC features. The same agreement between the BNN and DNN is also seen in Botometer features with the BNN obtaining an AUC of $0.973\pm 0.001$ and the DNN obtaining an AUC of $0.975$. In both cases, the RF outperforms the DNN and BNN due to its ability to operate more efficiently with smaller training sample sizes.
Figure 2: Overlayed ROC Curves: Receiver Operating Characteristic (ROC) Curves for the Bayesian Neural Network (BNN), Deep Neural Network (DNN) and Random Forest (RF), trained on BLOC features (left) and Botometer features (right). The uncertainty band on the BNN curves is obtained through a bootstrapping method, in which we sample the posterior over the weights to obtain uncertainties on the False Positive Rate (FPR) and True Positive Rate (TPR) at each threshold. Note the DNN and BNN perform consistently within error. RF outperforms the networks due to its increased ability to operate datasets with lower statistics more efficiently.
Figure 3: Uncertainty as a Function of Probability: Top row of plots are for the model trained on BLOC features, bottom row is for models trained on Botometer features. Note the parabolic like shape of the epistemic distribution, with maximum uncertainty around the decision boundary ($<p_{bot}>=0.5$).
We consolidated the uncertainty quantification by running a closure test, consisting of training the network (i) without including the aleatoric term, thereby predicting only the epistemic uncertainty, and (ii) including both aleatoric and epistemic terms. Details on how these two scenarios can be implemented are discussed in the Methods section and are also described in [38]. We demonstrated through a Z-score test that the epistemic uncertainties obtained on the accounts (considering the results from methods (i) and (ii) at the account level) produce consistent epistemic uncertainties (all values are within $|\text{Z}|\leq 0.5$). This closure test supports the fact that the quantified aleatoric uncertainty is decoupled from the epistemic uncertainty, with the latter appearing, on average, different from the former.
After consolidating our results, as discussed, we compare the performance of our network with respect to state-of-the-art works utilizing BLOC features [35] and Botometer [12] with random forest (RF) as a classifier, as done in those papers. We evaluate the performance using precision, recall, and F1 metrics. Our method outperforms other methods in terms of recall and F1, and is nearly on par in terms of precision, as shown in Table 1.
BLOC Botometer
Subset Model Precision Recall F1 Precision Recall F1
Human (Test Split) BNN 0.93 0.92 0.92 0.94 0.94 0.94
DNN 0.92 0.91 0.91 0.93 0.93 0.93
RF [39] 0.92 0.94 0.93 0.94 0.95 0.94
Bot (Test Split) BNN 0.92 0.93 0.93 0.94 0.94 0.94
DNN 0.91 0.92 0.91 0.93 0.93 0.93
RF [39] 0.94 0.91 0.93 0.95 0.94 0.94
Excess Human BNN 1.0 0.92 0.96 1.0 0.94 0.97
DNN 1.0 0.44 0.61 1.0 0.50 0.67
RF [39] 1.0 0.94 0.97 1.0 0.95 0.98
Table 1: Performance Comparison of Classifiers Using BLOC and Botometer Features: BNN used in our work, which is compared to a DNN and to a RF with features from BLOC and Botometer. Precision, recall and F1 have been computed for human accounts and bot accounts (using the test dataset which is a mixture of human and bot accounts), and using other bot accounts only not present in our test dataset to test generalization.

### v2

@@ Abstract
Keywords:
Uncertainty quantification, Bayesian neural network, Social media, Bot detection
journal: EPJ Data Science
1 Introduction
@@
Social media platforms have fundamentally transformed global communication, enabling the near-instant dissemination of information. Platforms like Facebook, YouTube, and Twitter/X, have over two billion monthly active users Dixon (2024) and empower individuals to broadcast their thoughts easily. However, the popularity that social media enjoys has incentivized malicious actors such as tech-savvy individuals or governments Twitter (2022), to deploy social bots to influence organic discourse and manipulate social media users for economic or political profit. Social bots Ferrara et al. (2016); Cresci (2020) — accounts controlled partly or fully by software — have been used to artificially boost the popularity of political candidates Ratkiewicz et al. (2011), spread conspiracy theories Lazer et al. (2018); Grinberg et al. (2019) during health crises Himelein-Wachowiak et al. (2021), and to manipulate the stock market Cresci et al. (2019); Pacheco et al. (2021).
@@
According to a 2023 study that analyzed 1 million tweets during a Republican debate and a Donald Trump interview, over 1,200 bot accounts were identified as spreading false narratives, highlighting the significant impact of bots during major events Graham and FitzGerald (2023). Anecdotal reports further underscore the prevalence of bot accounts, with some users suggesting that a substantial fraction of interactions on the platform are bot-generated. Despite enhanced verification processes, verified accounts continue to promote fraudulent schemes, illustrating the ongoing challenge in curbing bot activity.
@@
The issue of bot prevalence has also featured prominently in high-profile disputes, such as Elon Musk’s acquisition of Twitter. Musk’s legal team, using Botometer Yang et al. (2022), an online tool for identifying spam and fake accounts, claimed that 33% of “visible accounts” were “false or spam”. However, Botometer’s creator, Kaicheng Yang, criticized this claim and their methodology, stating that the figure was misleading and disclosed that Musk’s team had not consulted him before using the tool Clayton (2023). Perhaps even more concerning is the adaptation of Artificial Intelligence (AI) to bypass security systems deemed “prove you’re not a robot tests.” This concern has been echoed by Musk, Fig. 1 (left), as these tests are commonly used as initial filters for the prevention of inauthentic and/or malicious bot accounts. Moreover, accounts that successfully bypass such tests are increasingly becoming more “human-like,” as indicated by the dimensionality reduced representations of accounts in Fig. 1 (right). The figure illustrates three levels: the top represents the ground truth distribution, where many bots appear human-like. The middle plane shows classifier predictions, with blue favoring humans, red for bots, and white indicating indecision. The bottom plane depicts prediction uncertainty, with darker regions corresponding to higher uncertainty, aligning with indecisive areas in the middle plane. Musk has suggested that implementing a paid subscription model could be an effective way to combat bots on X (formerly Twitter). While the introduction of a paid subscription model aimed to enhance verification, concerns remain that it may not have fully addressed the issue of inauthentic accounts. Some argue that this approach could allow certain accounts to obtain verification status (blue check mark) through payment, potentially lending them an appearance of authenticity. This issue has been highlighted by the European Commission, which has claimed that Twitter was in violation of the Digital Services Act Commission (2024). Similarly, the decision to introduce a high-cost paywall for access to Twitter’s research Application Programming Interface (API) has posed challenges for researchers in studying and addressing bot activity.
@@
Figure 1: Impact of Bot Crisis and Uncertainty on Predictions (Our Work): As stated by Musk through X (left), the ability of bots to replicate human behavior and bypass security measures has increased dramatically with the advent of AI. Bot accounts are able to more efficiently mask themselves within the human population on social media platforms. This is shown through dimensionally reduced representations (right), in which we show three planes: (i) the true distributions, where we introduce an offset between human and bot points to ease visualization; (ii) the expected probability of an account being a bot as produced by our network, along with (iii) the associated uncertainty across the feature space, represented by the epistemic and aleatoric components added in quadrature. For (ii) and (iii) we use Gaussian process regression Rasmussen (2003) for visualization purposes. Uncertainty is greater in regions where ambiguity is higher and the two classes overlap.
@@
Given the serious harm that social bots wielded by bad actors, pose to democracy Schiffrin (2017); Woolley and Howard (2018), public health Tasnim et al. (2020); Allington et al. (2021); Pierri et al. (2022), or the the economy Max Fisher (2013), researchers have responded by developing a broad range of bot detection tools. Various Machine-Learning methods have been trained to detect social bots with a combination of features extracted from the social network structure, content/profile attributes, and temporal patterns Ferrara et al. (2016). Alternatively, all of these feature types are combined into a single model Lee et al. (2011); Davis et al. (2016); Varol et al. (2017); Gilani et al. (2017); Yang et al. (2019); Sayyadiharikandeh et al. (2020). Bot-detection algorithms often start by modeling the characteristics of accounts as the first step to distinguish bot from human-like behaviors Cresci (2020). Accounts may be represented using user profile information Yang et al. (2020), content Yardi et al. (2010); Cresci et al. (2017c, a), actions Mazza et al. (2019), social network Beskow and Carley (2018), and temporal signatures Mazza et al. (2019).
@@
Despite the progress made, bot detection remains a challenging, unsolved problem that is fraught with uncertainty due to the heterogeneity of bot behaviors, training data, and detection algorithms Cresci et al. (2023). Consequently, it is not surprising for detection models to disagree on whether to label the same account as bot or human, since they are often trained on different datasets that are sensitive to different subsets of the signals of automation. In light of this, it is paramount that bot detection models provide uncertainty weights alongside the probability estimate that an account is bot-controlled, as this could determine if one should trust a prediction. However, existing bot detection algorithms focus exclusively on detection.
@@
Our work addresses both bot detection and the quantification of uncertainty at the account level, an important distinctive feature of this research. This dual focus is crucial as it allows us to leverage additional information related to the quantified uncertainty of each prediction, thereby enhancing decision-making and improving the reliability of bot classifications. Specifically, our approach facilitates targeted interventions for bots where predictions are made with high confidence and suggests a cautious approach, e.g., gathering more data for bots where predictions are uncertain. Also, our method is agnostic to bot-detection algorithm, which enables the inclusion of uncertainty measurement into existing bot detection systems. As discussed earlier, research at the nexus of uncertainty quantification and bot detection is both timely and novel. In subsequent sections, we present our methodology that adeptly separates aleatoric and epistemic uncertainties. Aleatoric uncertainty, arising from the inherent randomness in our bot datasets, is characterized using multiple features from the Behavioral Language for Online Classification (BLOC) framework Nwala et al. (2023), or features from Botometer Yang et al. (2022). Conversely, epistemic uncertainty originates from the limitations of the predictive model used. Previous studies such as Taheri (2022) have employed Deep Ensemble and Stochastic Weight Averaging (SWA) for uncertainty estimation, yet these methods have notable limitations Abdar et al. (2021). SWA does not differentiate between the two types of uncertainties and requires supplementary techniques like dropout or bootstrapping for effective uncertainty quantification. Similarly, while Deep Ensembles capture epistemic uncertainty, they fail to distinctly separate it from aleatoric uncertainty.
@@
In contrast, our approach utilizes Bayesian methods, which provide a comprehensive and theoretically grounded framework for distinguishing and quantifying different types of uncertainties. We demonstrate that the computational demand of our Bayesian Neural Network (BNN) is negligible given the complexity of the problem at hand, and agnostic to input, e.g. BLOC or Botometer features. This approach positions our methodology as exceptionally capable in the critical domain of bot detection, offering a robust mechanism for managing uncertainty quantification.
In Sec. 2.1 we provide a literature review of bot detection methods, along with Uncertainty Quantification within the scope of bot detection in Sec. 2.2. In Sec. 3 we provide detailed descriptions of the feature extraction methods (BLOC and Botometer), and the Bayesian Deep Learning methods used. Sec. 4 details the experimental setup in terms of datasets, training and inference procedures, followed by Sec. 5 in which we present our results and performance metrics obtained through the implementation of these methods. Finally, Sec. 6 offers a summary of our findings and conclusions. The contributions of our work are as follows:
  • We introduce the first fully Bayesian Deep Learning method of uncertainty quantification in the space of bot detection; capable of providing both epistemic (model) and aleatoric (stochastic) uncertainties.
  • Our method provides uncertainty-aware decisions at the account level. Capable of improving performance of the architecture over baselines.
  • Our method is agnostic to bot detection features, implying usage in downstream tasks from other pre-processing schemes or Deep Learning based feature extraction algorithms.
2 Related Works
2.1 Bot Detection
Social media abusers utilize different tactics to manipulate their audiences for political Woolley and Howard (2018) or economic Max Fisher (2013) gain. One of of the earliest forms of abuse was spamming by spam bots (software-controlled accounts). Spam accounts were easy to detect because they lacked meaningful profile information and/or demonstrated naive behaviors Yardi et al. (2010); Lee et al. (2011). However, following the development of effective spam account detection methods (e.g., Markines et al. (2009); Jin et al. (2011); Zheng et al. (2015); Stringhini et al. (2010); Khan et al. (2015); Rastogi et al. (2020); Rastogi and Mehrotra (2017)), spam bots evolved to social bot accounts Ferrara et al. (2016); Cresci (2020). Similar to spam bots, social bots are controlled fully or partly by software, but are more sophisticated. For example, some accounts have detailed profiles, either stolen from real users or generated by deep neural networks Nightingale and Farid (2022). Some can interact with actual humans or mimic human behaviors by generating human-like content with ChatGPT Yang and Menczer (2024) and build social connections Cresci (2020). Others like cyborgs Chu et al. (2012, 2010) cycle between human and bot-like behaviors.
Various Machine-Learning methods have been trained to detect social bots with a combination of features extracted from the social network structure, content/profile attributes, and temporal patterns Ferrara et al. (2016). Alternatively, all of these feature types are combined into a single model Lee et al. (2011); Davis et al. (2016); Varol et al. (2017); Gilani et al. (2017); Yang et al. (2019); Sayyadiharikandeh et al. (2020); Cai et al. (2024). Bot-detection algorithms often start by modeling the characteristics of accounts as the first step to distinguish bot from human-like behaviors Cresci (2020). Accounts may be represented using user profile information Yang et al. (2020), content Yardi et al. (2010); Cresci et al. (2017c, a), actions Mazza et al. (2019), social network Beskow and Carley (2018), and temporal signatures Mazza et al. (2019).
The algorithms described here only produce probabilities estimating the likelihood that accounts are bots (software controlled). Since bot detection remains a challenging unsolved problem due to the heterogeneity of bot behaviors, training data, and detection algorithms Cresci et al. (2023), different detection methods could disagree on the label to assign to the same account. Consequently, an important novel contribution of this work is addressing both bot detection and the quantification of uncertainty at the account level. This enables us to leverage additional information related to the quantified uncertainty of each prediction, thereby enhancing decision-making and improving the reliability of bot classifications.
2.2 Uncertainty Quantification in Bot Detection
Uncertainty Quantification (UQ) is an increasingly critical component of decision-making, particularly with the adoption of AI-centric pipelines in the domain of bot detection. These pipelines often lack transparency due to their black-box nature and are typically unable to provide introspective confidence measures during inference. Consequently, UQ has emerged as a powerful tool, enabling fine-grained decision-making and identifying regions of low confidence (unreliability) within a model’s output space.
A natural approach to UQ involves Bayesian methods, which enable the decomposition of uncertainty into two distinct sources: epistemic and aleatoric. Epistemic uncertainty arises from limitations in the model itself, while aleatoric uncertainty reflects the inherent randomness in the data. Despite the critical importance of uncertainty-informed decision-making in managing interactions on social platforms, there has been limited exploration of UQ in the context of bot detection, particularly within the space of modern deep learning architectures.
Existing methods such as Stochastic Weight Averaging (SWA) Taheri (2022) are constrained by their reliance on approximating the posterior distribution through stochastic weight updates during the final training iterations. These approaches typically require additional techniques to produce meaningful uncertainty estimates. Similarly, Naive Bayes methods Gamallo and Almatarneh (2019); Kirn and Hinders (2021) often fall short due to their simplistic assumptions, rendering them inadequate for capturing both epistemic and aleatoric uncertainties.
In contrast, we present the first Bayesian deep learning framework for bot detection, which effectively separates and assesses both epistemic and aleatoric uncertainties. Our approach leverages state-of-the-art Bayesian deep learning techniques, inspired by advancements in Computer Vision (CV) Kendall and Gal (2017), alongside Multiplicative Normalizing Flows (MNF) Louizos and Welling (2017), to model complex posterior distributions. Our network design philosophy is inherited from those developed in Nuclear Physics Fanelli and Giroux (2024). A detailed description of the implemented methodology is provided in Sec. 3.2.
3 Methods
In Sec. 3.1 we describe the feature extraction methods used to form inputs to our bot detection pipelines, namely BLOC Nwala et al. (2023) and Botometer Yang et al. (2022). We will then describe the inner working of our Bayesian approach in Sec. 3.2.
3.1 Feature Extraction
BLOC
The BLOC framework Nwala et al. (2023) provides formal languages that represent the behaviors of social media accounts irrespective of social media platform, user-agent (human or bot), or intent (malicious or benign). The BLOC formal languages are defined by a set of alphabets (action and content) and rules for generating BLOC strings which are tokenized to produce BLOC words. BLOC words which represent the behaviors of social media accounts, consist of symbols drawn from distinct alphabets representing an account’s actions and content. Fig. 2a illustrates a possible representation of a sequence of tweets by three different Twitter accounts, @NASA, @Alice, and @Bob. The @NASA account replied to a tweet, posted a tweet, and then re-shared (retweet) a tweet, resulting in the BLOC action sequence: $p.T.r$. Here, each action (e.g., reply to) is represented by a single symbol (e.g., $p$), and the dots represent long pauses (e.g., $>$ 1 minute) between actions.
Once generated, BLOC strings are be tokenized (Fig. 2b) into bi-grams (two-letter words), then we can represent any social media account as a vector of BLOC words. A collection of point vectors corresponding to multiple accounts make up the BLOC matrix in Fig. 2c. In this vector space model, each account is represented as a point $(w_{1},w_{2},...,w_{k})$ in $k$-dimensional vector space where each dimension $i$ corresponds to a BLOC word. The weight $w_{i}$ represents how well an account is described by word $i$. For each account $a$, we instantiated $w_{i}$ with the TF-IDF weight Sparck Jones (1972), the product of the term frequency (TF) and the inverse document frequency (IDF):
$w_{i}(a)=f_{i}(a)\left(1+\log\frac{D}{d_{i}}\right)$ (1)
Botometer
Botometer,111botometer.org is a publicly available supervised machine learning system that classifies a given Twitter account as bot or human-controlled. Since the release of the first version in 2016, Botometer Davis et al. (2016) (formerly BotOrNot) has been cited over 2,000 times and used extensively by research published across diverse venues including Science Vosoughi et al. (2018) and Nature Pennycook et al. (2021). Like any Machine-Learning model, Botometer is not perfect, but with an F1 of 0.77 Sayyadiharikandeh et al. (2020) which was calculated from a heterogeneous dataset, it remains one of the most robust methods for bot detection.
Botometer-V4 (the current version of Botometer at the time of writing) Sayyadiharikandeh et al. (2020) utilizes over 1,000 features that can be grouped into six categories that focus on different account characteristics including, metadata from the accounts (e.g., numbers of friends and followers), retweet and mention networks, temporal features (e.g., frequency of posts), content information, and sentiment. In the deployed system, different classifiers in an ensemble are trained on different accounts types, and then these classifiers vote to obtain the final bot score Sayyadiharikandeh et al. (2020). Here however, we only utilize the representation power of Botometer by extracting its features which serve as input to BNN and DNN.
3.2 Bayesian Neural Networks
BNNs are extensions of traditional DNNs, aiming to optimize distributions of weights at each layer. [^2] In contrast to their deterministic counterparts, which define a fixed set of weights over the model. The aim of BNNs is to approximate a posterior distribution over the weights, given a dataset $q(\mathbb{W}|D)$. This allows predictions of quantities through a posterior distribution $q(\mathbb{y}|\mathbb{x},D)$, integrated over the weights. The formulation of such a posterior is intractable and therefore one must turn to Bayesian inference techniques during optimization. Traditional approaches define the posterior distribution to be a fully factorized Gaussian (diagonal covariance) $q(\mathbb{W})$, such that the evidence-lower bound (ELBO) can be maximized between the approximated posterior and the prior. Note that maximizing the ELBO is equivalent to minimizing the KL Divergence. This assumption, although more advantageous, tends to be limiting in terms of learned network complexity. Another approach provided in Louizos and Welling Louizos and Welling (2017), is to instead approximate the posterior as a product of fully factorized Gaussian and a mixing density, Eq. 2, where $q(\mathbb{z})$ is a vector of random variables. The random variables act multiplicatively on the means of the mixing density to reduce computational complexity.